# Driving Animatronic Robot Facial Expression From Speech

### v2

@@ Abstract
Animatronic robots aim to enable natural human-robot interaction through lifelike facial expressions. However, generating realistic, speech-synchronized robot expressions is challenging due to the complexities of facial biomechanics and responsive motion synthesis. This paper presents a principled, skinning-centric approach to drive animatronic robot facial expressions from speech. The proposed approach employs linear blend skinning (LBS) as the core representation to guide tightly integrated innovations in embodiment design and motion synthesis. LBS informs the actuation topology, enables human expression retargeting, and allows speech-driven facial motion generation. The proposed approach is capable of generating highly realistic, real-time facial expressions from speech on an animatronic face, significantly advancing robots’ ability to replicate nuanced human expressions for natural interaction.
@@
Accurately replicating human facial expressions is crucial for natural human-robot interaction [1, 2, 3]. Speech-synchronized, lifelike expressions enable social robots to achieve genuine emotional resonance with users [4, 5], underscoring the need to advance speech-driven expression generation for animatronic faces [6], heralding a paradigm shift in human-robot interactions.
@@
However, generating seamless, real-time animatronic facial expressions from speech presents two main challenges: (1) replicating the intricate biomechanics of human facial musculature [7, 8, 9, 10, 11, 12, 13, 14], and (2) generating nuanced human expressions through responsive algorithms based on advanced imitation learning [15, 16, 17, 18, 19]. Overcoming these challenges necessitates a comprehensive approach that integrates embodiment design and motion synthesis.
@@
Conventional muscle-centric embodiment design approaches attempt to replicate human anatomy, but cloning complex biomechanics presents significant engineering obstacles. Importantly, the primary objective is achieving realistic facial skinning motions rather than internal muscle movements. Consequently, this paper proposes a novel skinning-centric embodiment design approach.
@@
Current methods animate robot facial skinning through 3D landmark alignment between humans and robots [15, 17]. However, relying on sparse landmarks has inherent limitations: (1) insufficient capture of expression intricacies, resulting in oversimplification; (2) coupling of facial shapes and expressions, leading to inconsistencies from varying shapes; (3) topological differences hindering viable landmark motion transfer, limiting adaptability; and (4) confinement of edits to specific landmarks, restricting semantic adjustability. Thus, a 3D landmark-free skinning representation is required.
Figure 1: Expressive speech-driven facial dynamics achieved by the developed animatronic robot face. The figure showcases the realism and diversity of the generated robot facial expressions, synchronized with the corresponding speech input over time.
@@
To address these challenges, this paper proposes a principled approach leveraging linear blend skinning (LBS) for embodiment design and motion synthesis. For embodiment, LBS guides a skinning-oriented actuation topology optimized for blend skinning objectives while referencing facial anatomy [20]. Tendon-driven control points actuate the skin, enabling dynamic expressions with minimal spatial constraints. For synthesis, LBS facilitates motion retargeting from human demonstrations into robot skinning references. A speech-driven model is further proposed that generates highly realistic, lip-synchronized LBS-based skinning motions in real-time through imitation learning. Significantly, the LBS-based motions are semantically editable, allowing embodiment adaptation without retraining.
@@
Figure 2: The proposed approach for creating a speech-driven animatronic robot face using LBS. The approach comprises three major components: (1) skinning-oriented robot development designs and constructs the animatronic face paired with a kinematics simulator based on the target skinning appearance, (2) skinning motion imitation learning involves learning an LBS-based model from 3D human demonstrations to generate facial motions from speech, and (3) speech-driven robot orchestration generates animatronic facial expressions during inference by utilizing the developed platform, simulator, and learned model.
@@ II Related Works
Animatronic Robot Face: The evolution of animatronic robot faces can be categorized into two main phases: (1) early approaches that prioritized hardware design with pre-programmed expressions, and (2) recent studies that incorporate human motion transfer techniques. Pioneering works [7, 8, 9, 10, 11, 12] relied heavily on pre-programmed hardware and expressions, which severely limited their ability to generalize beyond a fixed set of postures. In recent years, researchers have made significant strides by integrating human motion transfer techniques, such as active appearance models [15], genetic algorithms [16], visual mimicry learning [17], Bayesian optimization [18], and MAP-Elites algorithms [19]. While these approaches have greatly enhanced the flexibility of animatronic robot faces, they remain fundamentally constrained by their reliance on mimicking observed human expressions through a master-slave mapping paradigm. Consequently, the generation of real-time, speech-driven facial expressions on animatronic platforms remains an unresolved challenge, impeding natural human-robot interaction.
Facial Expression Synthesis: Speech-driven facial expression synthesis has undergone significant advancements in generating photo-realistic talking head videos [21, 22, 23, 24]. These approaches map speech to target video domains, yielding impressive visual results. More pertinent to this work are methods for speech-driven 3D facial animation that control full vertex-level facial skinning [25, 26, 27, 28, 29]. Although these techniques excel at creating realistic virtual renditions, they do not address the distinct challenges inherent in physical robots. This work, while related, distinguishes itself by focusing on the generation of robot facial expressions from speech on real animatronic platforms. This task presents unique difficulties that remain largely unexplored, positioning this research at the forefront of this emerging field.
@@ III-A Approach Overview
Fig. 2 presents the proposed approach for creating a speech-driven animatronic robot face using LBS. It comprises three key components: First, skinning-oriented robot development designs and constructs the animatronic platform paired with a kinematics simulator based on the target skinning appearance. The LBS motion space is designed following blendshape design protocols [30] to enable seamless motion transfer. The actuation topology and simulator are developed concurrently to match this motion space, while the robot face is constructed considering physical constraints. Second, skinning motion imitation learning involves learning an LBS-based model from 3D human demonstrations to generate robot facial motions. Human blendshapes are carefully designed to align semantically with the robot blendshapes, ensuring consistent and expressive motion transfer. The learned model takes speech as input and outputs skinning motions as reference signals. Finally, during inference, speech-driven robot orchestration generates expressions on the animatronic face using the learned model and the developed simulator. Inverse kinematics is solved online to compute actuator commands from the generated skinning motions, enabling real-time, speech-driven expressions.
Figure 3: The proposed skinning-oriented robot design. The figure comprises two primary components: (1) LBS-oriented kinematics design to achieve actuation topology for the facial muscular system that matches the designed LBS motion space and references facial anatomy, and (2) electro-mechanical design and development accounting for physical constraints of the embodiment, including key mechanical components of the skin, skeleton and muscular system, as well as the electrical control system.
@@ III-B LBS Representation
LBS serves as the core representation in the proposed approach, offering several key advantages. First, it compacts intricate expressions into a small set of blendshape coefficients, facilitating human-like motions with minimal data. Second, as an expression-dependent but subject-invariant representation, LBS enables consistent performance across different faces by generating the same blendshape coefficients. This decouples embodiment design from motion synthesis, making the LBS-based motion synthesis model generalizable across diverse animatronic embodiments and semantically re-editable before retargeting, allowing adaptation without retraining.
@@
where $\mathbf{T}^{R}\in\mathbb{R}^{3U}$ is the robot face template with $U$ vertices, and $B_{E}^{R}(\boldsymbol{\theta};\mathbf{\mathcal{E}}^{R}):\mathbb{R}^{\boldsymbol{\theta}}\mapsto\mathbb{R}^{3U}$ is the expression skinning function with $B$ robot blendshape bases $\mathbf{\mathcal{E}}^{R}$ and coefficients $\boldsymbol{\theta}$.
To enable motion transfer, human blendshapes $\mathbf{\mathcal{E}}^{H}$ are designed to align semantically with $\mathbf{\mathcal{E}}^{R}$. The human mesh topology matches that of the captured 3D human demonstrations, with the mean face shape of demonstration subjects serving as the human face template $\mathbf{T}^{H}$. The human facial skinning function $T^{H}(\boldsymbol{\theta}):\mathbb{R}^{\boldsymbol{\theta}}\mapsto\mathbb{R}^{3V}$ is
@@
For the blendshape design, the Apple ARKit standard (excluding tongue-out) is adopted for its semantic meaningfulness, comprehensiveness, and interoperability.
@@ IV-A LBS-Oriented Kinematics Design
The primary objective in designing the facial muscular system is to reproduce the target LBS-based motion space rather than precisely replicating the intricate human musculature. An acceptable system only needs to match the desired skinning motions, not anatomical accuracy.
@@
Through iterative optimization, referencing LBS objectives and facial anatomy, the kinematics is designed. As depicted in Fig. 3, it contains a total of $21$ control points: $4$ for eyebrows, $4$ for eyelids, $2$ for eyeballs, $2$ for the nose, $2$ for cheeks, $6$ for the mouth, and $1$ for the jaw. The color-coded points show the control point locations, and lines indicate the B6DOF motion bounds. This optimized design strikes a balance between motion flexibility and system complexity.
@@ IV-B Electro-Mechanical Design and Development
Figure 4: The proposed speech-driven facial skinning motion imitation learning method. The model (blue block) is trained through the train branch (red block) to generate blendshape coefficients from input speech by imitating human facial skinning motions. During inference (orange block), the model predicts blendshape coefficients, which are further decoded by the robot LBS decoder to generate robot facial skinning motions as reference signals for the kinematics simulator.
Mechanical Design: The primary objectives in the mechanical design are physically realizing the facial muscular motion space defined by the control points while accounting for physical constraints. To achieve this, a tendon-driven actuation approach is proposed, inspired by human facial anatomy. This approach enables the remote location of actuators, providing power under spatial limitations.
@@
The skeletal and muscular system design, shown in Fig. 3, is carefully crafted to constrain the control points’ B6DOF motions, ensuring they closely match the target LBS-based motion space. Although the facial skinning is simulated, it serves as the ultimate output of the system. To achieve the desired level of realism, a silicone mask was iteratively designed for the facial skin, taking into account factors such as shape, hardness, and thickness, inspired by human skin. The mask’s soft silicone provides exceptional flexibility, allowing the robot to achieve the desired motions with a high degree of fidelity. Experimental results validate the successful achievement of the designed robot blendshapes, serving as a compelling indicator of the effectiveness of the proposed skinning-oriented design approach.
Electrical Design: The primary objective in the electrical design is to enable real-time skinning motion tracking by the animatronic face. To achieve responsive $25\mathrm{Hz}$ control for robot orchestration, the system must actuate each control point across its full motion range within $40\mathrm{ms}$.
To meet these stringent real-time requirements, the design leverages specific components for signal processing and actuation. First, the ESP32 microcontroller is chosen for its efficient on-chip signal processing capabilities. Next, two PCA9685 boards provide interfaces for driving the total $31$ servo motors required for articulation. In particular, $3$ high-torque DS3115 servos are selected for the neck module to enable rapid rotations under the payload of the robot head. The $28$ lower-torque LFD-01M servos optimize cost while still providing adequate speed for the finer facial motions. With a $0.12\mathrm{sec}/60^{\circ}$ at $4.8\mathrm{V}$, the servos exceed the $40\mathrm{ms}$ actuation constraints when powered by a $5\mathrm{V}$ lithium battery.
The high-level system uses a GPU-accelerated PC, which generates motion references from the learned model and runs the simulator to solve inverse kinematics to generate servo commands for the ESP32 microcontroller. Experimental results validate accurate real-time tracking, demonstrating that the proposed electrical design succeeds in enabling dynamic facial expression control on the developed robot platform.

### v3

@@ Abstract
Animatronic robots hold the promise of enabling natural human-robot interaction through lifelike facial expressions. However, generating realistic, speech-synchronized robot expressions poses significant challenges due to the complexities of facial biomechanics and the need for responsive motion synthesis. This paper introduces a novel, skinning-centric approach to drive animatronic robot facial expressions from speech input. At its core, the proposed approach employs linear blend skinning (LBS) as a unifying representation, guiding innovations in both embodiment design and motion synthesis. LBS informs the actuation topology, facilitates human expression retargeting, and enables efficient speech-driven facial motion generation. This approach demonstrates the capability to produce highly realistic facial expressions on an animatronic face in real-time at over 4000 fps on a single Nvidia RTX 4090, significantly advancing robots’ ability to replicate nuanced human expressions for natural interaction. To foster further research and development in this field, the code has been made publicly available at: https://github.com/library87/OpenRoboExp.
@@
Accurately replicating human facial expressions is crucial for natural human-robot interaction [1, 2, 3]. Recent studies have incorporated human motion transfer techniques for animatronic faces [4, 5, 6, 7, 8], primarily focusing on mimicking observed human expressions through leader-follower mapping. However, to achieve genuine emotional resonance with humans, social robots require speech-synchronized, lifelike expressions [9, 10]. This necessitates a shift towards speech-driven approaches for generating dynamic and contextually appropriate facial expressions. Such advancements have significant implications for various applications, such as entertainment, education and healthcare, where natural and expressive human-robot interaction is paramount.
This paper introduces the first principled approach for creating an animatronic robot face capable of generating expressions directly from speech, marking a significant advancement in creating dynamically expressive robot faces and enhancing the potential for natural human-robot interaction. Fig. 1 showcases the realism and diversity of the generated robot facial expressions, synchronized with the corresponding speech input over time.
@@
Generating seamless, real-time animatronic facial expressions from speech presents two main challenges: (1) replicating the intricate biomechanics of human facial musculature [11, 12, 13, 14, 15, 16, 17, 18], and (2) generating nuanced human expressions through responsive algorithms based on advanced imitation learning [4, 5, 6, 7, 8]. Overcoming these challenges necessitates a comprehensive approach that integrates embodiment design and motion synthesis.
@@
Conventional muscle-centric embodiment design approaches attempt to replicate human anatomy, but face significant engineering obstacles. These include the complexity of mimicking the multitude of facial muscles, their interconnections, and their subtle interactions. Additionally, the miniaturization of actuators to fit within the confined space of a robot face while maintaining the required force and precision presents considerable challenges. The key insight of this research is that achieving realistic facial skinning motions, rather than replicating internal muscle movements, is the primary objective. Consequently, this paper proposes a novel skinning-centric embodiment design approach.
@@
Current methods animate robot facial skinning through 3D landmark alignment between humans and robots [4, 6]. However, this sparse landmark approach has inherent limitations: (1) insufficient capture of expression intricacies, resulting in oversimplification; (2) coupling of facial shapes and expressions, leading to inconsistencies from varying shapes; (3) topological differences hindering viable landmark motion transfer, limiting adaptability; and (4) confinement of edits to specific landmarks, restricting semantic adjustability. These limitations underscore the need for a 3D landmark-free skinning representation.
Figure 1: Dynamic animatronic robot facial expressions generated from speech. The figure shows the system’s capability to produce diverse and lifelike facial expressions in real-time, synchronized with the corresponding audio speech input. The waveform at the top represents the audio input, while the series of images below showcase the robot’s facial responses at different time points.
@@
To address these challenges, this paper proposes a principled approach leveraging linear blend skinning (LBS) for both embodiment design and motion synthesis. LBS enables efficient and controllable facial expressions by linearly combining predefined shape variations (blendshapes) from a neutral shape. For embodiment, LBS guides a skinning-oriented actuation topology optimized for blend skinning objectives while referencing facial anatomy. For synthesis, LBS facilitates motion retargeting from human demonstrations into robot skinning references. A speech-driven model is further proposed that generates highly realistic, lip-synchronized LBS-based skinning motions in real-time through imitation learning. This approach significantly advances upon recent works by offering enhanced expressiveness through continuous dense deformation, improved consistency and adaptability to different facial topologies, semantic editability without need to retrain for different robot faces, and efficient real-time performance crucial for natural human-robot interaction.
@@
The rest of the paper is organized as follows: Section II reviews related works in animatronic robot faces and facial expression synthesis. Section III introduces the proposed LBS-based approach. Section IV details the skinning-oriented robot development process. Section V elaborates on the skinning motion imitation learning method. Section VI presents experimental results, and Section VII concludes with a discussion of implications and future research directions.
@@ II Related Works
Figure 2: The proposed approach for creating a speech-driven animatronic robot face using LBS. The approach comprises three major components: (1) skinning-oriented robot development designs and constructs the animatronic face paired with a kinematics simulator based on the target skinning appearance, (2) skinning motion imitation learning involves training an LBS-based model from 3D human demonstrations to generate facial expressions from speech input, and (3) speech-driven robot orchestration generates animatronic facial expressions during inference by utilizing the developed platform, simulator, and learned model. The diagram highlights key development steps, outputs, and inference processes, demonstrating the end-to-end workflow from concept to final animatable robot face.
Animatronic Robot Face: The evolution of animatronic robot faces can be categorized into two main phases: (1) early approaches that prioritized hardware design with pre-programmed expressions, and (2) recent studies that incorporate human motion transfer techniques. Pioneering works [11, 12, 13, 14, 15, 16] relied heavily on pre-programmed hardware and expressions, which severely limited their ability to generalize beyond a fixed set of postures. In recent years, researchers have made significant strides by integrating human motion transfer techniques, such as active appearance models [4], genetic algorithms [5], visual mimicry learning [6], Bayesian optimization [7], and MAP-Elites algorithms [8]. While these approaches have enhanced the flexibility of animatronic robot faces, they remain fundamentally constrained by their reliance on mimicking observed human expressions through a leader-follower mapping paradigm. The creation of real-time facial expressions directly from speech on animatronic platforms remains an unexplored frontier, impeding natural human-robot interaction.
Facial Expression Synthesis: In parallel, speech-driven facial expression synthesis has seen substantial advancements in generating photo-realistic talking head videos [19, 20, 21, 22]. These approaches map speech to target video domains, yielding impressive visual results. More pertinent to this work are methods for speech-driven 3D facial animation that control full vertex-level facial skinning [23, 24, 25, 26, 27]. However, while these methods excel at creating realistic virtual renditions, they do not address the distinct challenges inherent in physical robots, such as actuation limitations, real-time control requirements, and physical constraints. Their direct applicability to animatronic platforms remains limited.
This work distinguishes itself by introducing the first principled approach for generating robot facial expressions directly from speech on real animatronic platforms. It bridges the gap between speech input and physical actuation, a critical step towards creating dynamically expressive robot faces capable of natural, speech-synchronized interactions. By addressing the unique challenges of translating speech into physical facial movements, this research opens new avenues for enhancing the expressiveness and naturalness of human-robot interaction.
@@ III-A Approach Overview
Fig. 2 presents the proposed approach for creating a speech-driven animatronic robot face using LBS. It comprises three key components: First, skinning-oriented robot development designs and constructs the animatronic platform paired with a kinematics simulator based on the target skinning appearance. The LBS motion space is designed following blendshape design protocols [28] (predefined bases of facial expressions) to enable seamless motion transfer. The actuation topology and simulator are developed concurrently to match this motion space, while the robot face is constructed considering physical constraints. Second, skinning motion imitation learning involves learning an LBS-based model from 3D human demonstrations to generate robot facial motions. Human blendshapes are carefully designed to align semantically with the robot blendshapes, ensuring consistent and expressive motion transfer. The learned model takes speech as input and outputs skinning motions as reference signals. Finally, during inference, speech-driven robot orchestration generates expressions on the animatronic face using the learned model and the developed simulator. Inverse kinematics is solved online to compute actuator commands from the generated skinning motions, enabling real-time, speech-driven expressions.
@@ III-B LBS Representation
The proposed approach leverages LBS, a widely-used technique in computer graphics for deforming 3D meshes. LBS represents a mesh as a weighted combination of a neutral shape and a set of predefined shape variations, called blendshapes. It offers key advantages: First, it compactly encodes intricate expressions into a small set of blendshape coefficients, facilitating human-like motions with minimal data. Second, as an expression-dependent but subject-invariant representation, LBS enables consistent performance across different faces by generating the same blendshape coefficients for the same expressions. This decoupling of embodiment design from motion synthesis allows the LBS-based motion synthesis model to generalize across diverse embodiments and enables semantic editability before retargeting without retraining.
Crucially, the approach relies on semantically aligning the design of human and robot blendshapes to enable motion transfer. By carefully designing these aligned blendshapes to represent the same set of facial expressions (e.g., eye blink, mouth smile), it ensures that the same blendshape coefficients, when applied to both human and robot, produce semantically corresponding expressions, even with differing underlying facial structures.
@@
where $\mathbf{T}^{R}\in\mathbb{R}^{3U}$ is the neutral (zero-pose) robot face with $U$ vertices, and $B_{E}^{R}(\boldsymbol{\theta};\mathbf{\mathcal{E}}^{R}):\mathbb{R}^{\boldsymbol{\theta}}\mapsto\mathbb{R}^{3U}$ is the expression skinning function with $B$ robot blendshape bases $\mathbf{\mathcal{E}}^{R}$ and coefficients $\boldsymbol{\theta}$. The expression skinning function $B_{E}^{R}$ linearly combines the robot blendshape bases $\mathbf{\mathcal{E}}^{R}$ using the coefficients $\boldsymbol{\theta}$ to generate the final facial deformation, allowing efficient representation and control of diverse expressions by adjusting the blendshape coefficients.
To enable motion transfer, human blendshapes $\mathbf{\mathcal{E}}^{H}$ are designed to align semantically with $\mathbf{\mathcal{E}}^{R}$. The human mesh topology matches that of the captured 3D human demonstrations, with the mean face shape of demonstration subjects serving as the neutral human face $\mathbf{T}^{H}$. The human facial skinning function $T^{H}(\boldsymbol{\theta}):\mathbb{R}^{\boldsymbol{\theta}}\mapsto\mathbb{R}^{3V}$ is
@@
For the blendshape design, the Apple ARKit standard (excluding tongue-out) is adopted for its semantic meaningfulness and comprehensiveness, enhancing practical interoperability by leveraging existing frameworks.
Figure 3: The proposed skinning-oriented robot design. The figure comprises two primary components: (1) LBS-oriented kinematics design, which showcases the facial mesh model with strategically placed control points for various facial features to achieve actuation topology for the facial muscular system that matches the designed LBS motion space and references facial anatomy, and (2) electro-mechanical design and development accounting for physical constraints of the embodiment, including key mechanical components of the skin, skeleton and muscular system, as well as the electrical control system. This comprehensive view demonstrates how the theoretical LBS model is translated into a functional, physically embodied animatronic face.
@@ IV-A LBS-Oriented Kinematics Design
The primary objective in designing the facial kinematics system is to reproduce the target LBS-based motion space, prioritizing functional equivalence over anatomical replication. A baby-like face serves as the target animatronic robot appearance, demonstrating the proposed skinning-centric approach’s applicability to various facial morphologies.
@@
As depicted in Fig. 3, the optimized kinematics design contains a total of $21$ control points: $4$ for eyebrows, $4$ for eyelids, $2$ for eyeballs, $2$ for the nose, $2$ for cheeks, $6$ for the mouth, and $1$ for the jaw. The color-coded points show the control point locations, and lines indicate the B6DOF motion bounds. This design strikes a balance between motion flexibility and system complexity.
@@ IV-B Electro-Mechanical Design and Development
Mechanical Design: The mechanical design aims to physically realize the facial muscular motion space defined by the control points while accounting for physical constraints. A tendon-driven actuation approach, inspired by human facial anatomy, enables the remote location of actuators, providing power under spatial limitations.
@@
The skeletal and muscular system design, shown in Fig. 3, carefully constrains the control points’ B6DOF motions to match the target LBS-based motion space. This tendon-driven approach allows for complex, coordinated movements that closely mimic the desired facial expressions. A silicone mask, iteratively designed with considerations for shape, hardness, and thickness, serves as the facial skin. The mask’s soft silicone provides exceptional flexibility, enabling the robot to achieve desired motions with high fidelity, effectively translating the LBS-based design into physical form.
Figure 4: The proposed speech-driven facial skinning motion imitation learning method. The model architecture (blue section) comprises three key components: (1) a frame-level speech encoder that processes audio input and generates phoneme logits, (2) a speaking style encoder that captures individual speaking styles, and (3) an LBS encoder that generates blendshape coefficients. During training (red section), the model learns to imitate human facial skinning motions by minimizing the difference between generated and target expressions. In the inference branch (orange section), the trained model generates blendshape coefficients for the robot LBS decoder, producing robot-specific facial skinning motions as reference signals for the downstream kinematics simulator.
Electrical Design: The electrical design aims to enable real-time skinning motion tracking by the animatronic face. To achieve responsive $25\mathrm{Hz}$ control for robot orchestration, the system must actuate each control point across its full motion range within $40\mathrm{ms}$.
To meet these stringent real-time requirements, specific components are carefully selected based on their capabilities and performance characteristics. The ESP32 microcontroller is chosen for its efficient on-chip signal processing capabilities, enabling real-time control. Two PCA9685 boards provide interfaces for driving the total $31$ servo motors required for articulation. Three high-torque DS3115 servos are selected for the neck module due to their ability to enable rapid rotations under the payload of the robot head, while $28$ lower-torque LFD-01M servos optimize cost while still providing adequate speed for the finer facial motions. With a $0.12\mathrm{sec}/60^{\circ}$ at $4.8\mathrm{V}$, the servos exceed the $40\mathrm{ms}$ actuation constraints when powered by a $5\mathrm{V}$ lithium battery.
The high-level system uses a GPU-accelerated PC to generate motion references from the learned model and run the simulator to solve inverse kinematics, generating servo commands for the ESP32 microcontroller. This setup ensures that the complex LBS-based motions are accurately translated into physical movements in real-time.